The radius of the circle is 6 cm. The perpendicular
distance from the centre of the circle to the chord,
which is 8 cm in length, is .... ?

(A) 2.5 cm
(B) 13 cm
(C) 5 cm
(D) 17 cm​

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer:

we know that perpendicular drawn from the center to the chord divides the chord into two equal parts

so, AM=MB

=AB/2

=8/2

=4cm

using pythogoras therom in the ∆AMO

OM^2= AO^2-AM^2

= 6^2-4^2

=√20

=2√5

=5

Thanks